OpenStudy (anonymous):

6b greater than or equal to 3 over 5

OpenStudy (anonymous):

\[6b \ge 3/5\] \[b \ge \left( 3\over5 \right)\left( 1\over6 \right)\] \[b \ge \left( 1\over10 \right)\]

OpenStudy (anonymous):

number line?

OpenStudy (anonymous):

graph*

OpenStudy (anonymous):

money gang